A 0.075kg arrow is fire horizontally. The bowstring exerts an average force of 65N on the arrow over a distance of 0.09m. With what speed did the arrow leave on bow. Show complete solution.


1
Expert's answer
2021-12-20T10:31:40-0500

Find the acceleration:


"a=F\/m."

The speed:


"v=\sqrt{2ad}=\sqrt{2\frac Fmd}=\sqrt{2·\frac{65}{0.075}·0.09}=12.5\text{ m/s}."